Find the area of an isosceles triangle with legs of 12cm and a base of 6cm.
Answer by nerdybill(7384)

Find the area of an isosceles triangle with legs of 12cm and a base of 6cm.
.
area of any triangle = (1/2)bh
where
b is base
h is height
.
the problem gives you the base (b) as 6 cm.
You need to find the height (h). The height (h) splits the isosceles triangle in half and it is a right triangle -- use Pythagorean theorem to find h:
3^2 + h^2 = 12^2
9 + h^2 = 144
h^2 = 135
h = sqrt(135)
h = sqrt(5*3*3*3)
h = 9*sqrt(15)
.
area = (1/2)bh
area = (1/2)(6)(9*sqrt(15))
area = (3)(9*sqrt(15))
area = 27*sqrt(15) sq cm
area = 104.57 sq cm